OpenStudy (spyro517):

Line AB contains points A (8, −4) and B (1, −5). The slope of line AB is −7 negative 1 over 7 1 over 7 7

OpenStudy (anonymous):

m=1/7

OpenStudy (tkhunny):

Have you considered calculating it? \(\dfrac{Change\;in\;Y}{Change\;in\;X}\) Go!

OpenStudy (tkhunny):

Chris received a warning for that inappropriate response. Please SHOW YOUR WORK!

OpenStudy (tkhunny):

\(\dfrac{-5-(-4)}{1-8} = \dfrac{-1}{-7} = \dfrac{1}{7}\) Please read the Code of Conduct. http://openstudy.com/code-of-conduct Your declaration is in direct violation of the Terms and Conditions you agreed to when you registered.